# Cycles in spherical Deligne complexes and application to $K(\pi,1)$-conjecture for Artin groups

### v1

@@ Abstract
We show the $K(\pi,1)$-conjecture holds for Artin groups whose Dynkin diagrams are complete bipartite (edge labels are allowed to be arbitrary), answering a question of J. McCammond. Along the way, we treat several related families of hyperbolic type Artin groups, namely the $K(\pi,1)$-conjecture holds for all 3-dimensional hyperbolic type Artin groups, except one single example with Dynkin diagram $[3,5,3]$; and the conjecture holds for all quasi-Lannér hyperbolic type Artin groups up to dimension 4. We also treat several higher dimensional families.
Most of the article is about developing new methods of understanding combinatorial minimal fillings of certain types of cycles in spherical Deligne complexes or relative Artin complexes, via non-positive curvature geometry. Then we combine this with an approached to the $K(\pi,1)$-conjecture introduced by a previous article of the author to settle new cases of $K(\pi,1)$-conjecture.
In the appendix we list some related open questions and conjectures.
@@ 1 Introduction
1.1 Background and overview
Let $\{H_{\lambda}\}_{\lambda\in\Lambda}$ be a locally finite collection of affine complex hyperplanes in $\mathbb{C}^{n}$. The associated hyperplane arrangement complement is defined to the manifold $\mathbb{C}^{n}-\cup_{\lambda\in\Lambda}H_{\lambda}$. Despite the simple definition of such spaces and the long history of studying them, some fundamental aspects of their topology remain elusive; even many basic questions about their fundamental groups are widely open, except for a number of special cases.
One important scenario where these arrangement complements become comparably more tractable (though still mysterious) is that the collection of hyperplanes has extra symmetry – there is a group acting on $\mathbb{C}^{n}$ permuting the hyperplanes. As a basic example, we take the family of hyperplanes to be $\{z_{i}=z_{j}\}_{1\leq i\neq j\leq n}$ in $\mathbb{C}^{n}$. The symmetry group $S_{n}$ acts on $\mathbb{C}^{n}$ by permuting the coordinates; hence, it also permutes these hyperplanes. The associated arrangement complement is the configuration space of ordered $n$ points in $\mathbb{C}$, whose fundamental group is the pure braid group on $n$-strands.
The action of the symmetry group in this example is generated by reflections, and the chosen family of hyperplanes are reflection hyperplanes under such action. This has a vast generalization to any abstract reflection groups acting on $\mathbb{C}^{n}$, which gives rise to much more intricate collections of reflection hyperplanes.
Let $\Gamma$ be a finite simplicial graph with each edge labeled by an integer $\geq 2$ with its vertex set $S$. The Artin group (or Artin-Tits group) with presentation graph $\Gamma$, denoted $A_{\Gamma}$ or $A_{S}$, is a group whose generating set is $S$, and there is a relation of the form $aba\cdots=bab\cdots$ with both sides being alternating words of length $m$ whenever two vertices $a$ and $b$ are connected by an edge labeled by $m_{ab}$. The Coxeter group with presentation graph $\Gamma$, denoted $W_{\Gamma}$ or $W_{S}$, has the same generating sets and the same relators as the Artin group, with extra relations $v^{2}=1$ for each vertex $v\in\Gamma$.
A reflection of $W_{S}$ is a conjugate an element in $S$. Let $R$ be the collection of all reflections in $W_{S}$. Recall that $W_{S}$ admits a canonical representation $\rho:W\to GL(n,\mathbb{R})$, such that each element in $R$ acts as a linear reflection on $\mathbb{R}^{n}$, and the action $W_{S}\curvearrowright\mathbb{R}^{n}$ stabilizes an open convex cone $I\subset\mathbb{R}^{n}$, called the Tits cone, where the action of $W_{S}$ is properly discontinuous. For each reflection $r\in W_{S}$, let $H_{r}$ be the set of fix points of $\rho(r)$ in $I$. The collection of all such $H_{r}$ form an arrangement of hyperplanes in $I$. Define
@@ 1.1 Background and overview
The $K(\pi,1)$-conjecture for reflection arrangement complements, due to Arnol’d, Brieskorn, Pham, and Thom, predicts that the space $M(W_{S})$ is aspherical for any Coxeter group $W_{S}$. There is an induced action of $W_{S}$ on $M(W_{S})$, which is free and proper discontinuous. The quotient has its fundamental group isomorphic to the Artin group $A_{S}$.
@@
Historically, the interests in these topological spaces (i.e., reflection arrangement complements) come from connection with singularity theory: if $W$ is of type $A$, $D$, or $E$, then $M(W_{S})/W_{S}$ is the complement of the discriminant of the semi-universal deformation of a simple singularity of the same type [Bri70]. More recently, it is known that many of these topological spaces are closely related to  spaces of stability conditions on suitable triangulated categories [Bri09, Ike14, QW18, AW22, Hen22, DHL23, QZ23], see also the survey article [Hen24]. In a few sporadic cases, these topological spaces also arise as components of certain strata of the moduli space of abelian differential in genus 3 [LM12].
The study of the $K(\pi,1)$ conjecture is closely related to the study of Artin groups. Basic questions of Artin groups remain widely open [GP12a], despite their simple presentation.
To date, the $K(\pi,1)$-conjecture is largely open. It is known that the $K(\pi,1)$-conjecture for an arbitrary Artin group reduces to the $K(\pi,1)$-conjecture for Artin groups with complete presentation graph by the work of Charney-Davis, Ellis-Sköldberg and Godelle-Paris [CD95a, ES10, GP12b]. Among Artin groups with a complete presentation graph, the $K(\pi,1)$-conjecture is known in the 2-dimension case by Charney-Davis [CD95a], and a few families that are not exactly 2-dimensional but have some low dimensional features by work of Charney, Juhasz and Goldman [Cha04, Juh18, Gol22, Juh23]. Known cases of $K(\pi,1)$-conjecture in higher dimensions are sparse - Artin groups associated with finite reflection groups by Deligne [Del72], Artin groups associated with Euclidean reflection groups by Paolini-Salvetti [PS21] (based on previous work of McCammond-Sulway [MS17]), and a few examples of hyperbolic type Artin groups by Haettel and the author [HH23].
In a previous article [Hua23], we proved the $K(\pi,1)$-conjecture for new families of higher dimensional Artin groups by establishing a connection between the $K(\pi,1)$-conjecture and the properties of short cycles in the 1-skeleton of Deligne complexes of spherical Artin groups. This leads to the study of short cycles in spherical Deligne complexes, which is a major challenge in implementing the strategy in [Hua23] for more general Artin groups.
In this article and [Hua24], we develop new geometric and topological methods to understand certain types of cycles in the spherical Deligne complexes; as a consequence, we settle the $K(\pi,1)$-conjecture of several important classes of higher dimensional Artin groups. This article can be read independently, assuming a few facts from [Hua23].
The study of cycles in Deligne complexes is closely related to solving certain types of equations over Artin groups. This article focuses on using non-positive curvature geometry to understand these equations, via constructing new actions of certain Artin groups and some of their quotients on injective metric spaces, CAT$(0)$ spaces, spaces with convex geodesic bicombing, Helly graphs and weakly modular graphs. The companion article [Hua24] explores other methods to understand these equations, including combinatorics of arcs on surfaces, as well as solving equations in free groups.
Some ingredients here will be used in an upcoming article to treat the $K(\pi,1)$ problem of some complex hyperplane complements outside the realm of Artin groups.
1.2 Main results
As the $K(\pi,1)$-conjecture for general Artin groups is reduced to Artin groups with complete presentation graphs, our results will be mostly about Artin groups with complete presentation graphs. It is more convenient to use a different graph in this case. Given an Artin group with presentation graph $\Gamma$, its Dynkin diagram $\Lambda$ is obtained from $\Gamma$ by removing all open edges of $\Gamma$ labeled by $2$, and adding extra edges labeled by $\infty$ between vertices of $\Gamma$ that are not adjacent. Thus, two vertices of the Dynkin diagram $\Lambda$ commute if they are not adjacent. For example, the presentation graph of any braid group is a complete graph, but its Dynkin diagram is a line, which is more informative.
@@ 1.2 Main results
(=Theorem 10.8) Let $A_{S}$ be an Artin group such that its Dynkin diagram is complete bipartite. Then $A_{S}$ satisfies the $K(\pi,1)$-conjecture.
A complete bipartite graph is a join of two discrete sets. Artin groups with complete bipartite Dynkin diagram were used by McCammond [McC17, Section 8] to demonstrate how little we understand about Artin groups. When the Dynkin diagram is the star $K_{1,n}$ (the join of one vertex and $n$-vertices) and all edges are labeled by $3$, the only previously known cases of $K(\pi,1)$-conjecture were $K_{1,1}$ (type $A_{2}$) and $K_{1,2}$ (type $A_{3}$) in [FN62], $K_{1,3}$ (type $D_{4}$) in [Bri06, Del72] and $K_{1,4}$ (type $\widetilde{D}_{4}$) very recently in [PS21]. The $K(\pi,1)$-conjecture for the next diagram $K_{1,5}$, whose Coxeter group acts on the hyperbolic space $\mathbb{H}^{5}$ with finite volume fundamental domain, was already not known before. The edge labeling in Theorem 1.1 is allowed to be arbitrary.
The choice of the family with complete bipartite Dynkin diagrams is not arbitrary, as it is closely related (in a perhaps less obvious way) to several other natural families of Artin groups. As the $K(\pi,1)$-conjecture is known for Artin groups associated with spherical and Euclidean reflection groups [Del72, PS21], it is natural to ask the same question for hyperbolic type Artin groups, i.e., Artin groups associated with Coxeter groups acting on $\mathbb{H}^{n}$ with finite volume fundamental domain. This was completely understood in the case of $\mathbb{H}^{2}$ [CD95a], with an alternative proof [DPS22]. However, even in $\mathbb{H}^{3}$, we only know a few sporadic examples of Artin groups satisfying $K(\pi,1)$-conjecture [Cha04, HH23, Hua23].
En route to proving Theorem 1.1, we treat the $K(\pi,1)$-conjecture for all 3-dimensional hyperbolic type Artin groups, except a single example.
@@
(=Corollary 9.3) Let $W_{\Gamma}$ be a reflection group acting properly on $\mathbb{H}^{3}$ by isometries, with a finite volume fundamental domain. Assume $W_{\Gamma}$ is not $[3,5,3]$. Then the $K(\pi,1)$-conjecture holds true for the associated Artin group $A_{\Gamma}$.
Here $[3,5,3]$ denotes the Coxeter group whose Dynkin diagram is a linear graph with three edges, such that consecutive edges are labeled $3,5,3$.
For general hyperbolic type Artin groups, the most fundamental classes are the Lannér and quasi-Lannér groups. A hyperbolic type Artin group is Lannér if the action of the associated Coxeter group has a fundamental domain being a compact simplex, and is quasi-Lannér if the associated Coxeter group has a fundamental domain being a non-compact finite volume simplex. A Dynkin diagram is Lannér (or affine or quasi-Lannér) if the associated Coxeter/Artin group satisfies the same property. We use $A_{\Lambda}$ to denote the Artin group with Dynkin diagram $\Lambda$.
@@
(=Corollary 11.4) Let $A_{\Lambda}$ be a quasi-Lannér hyperbolic type Artin group of dimension $\leq 4$. Then the $K(\pi,1)$-conjecture holds for $A_{\Lambda}$.
Theorem 1.3 is a consequence of a more general theorem treating higher dimensional Artin groups, which we will explain now. A key feature of the Dynkin diagram $\Lambda$ of a quasi-Lannér group, is that $\Lambda$ has a core, corresponding to the cusps of the action on $\mathbb{H}^{n}$, which is an affine subdiagram $\Lambda^{\prime}$, such that for any vertex $s\in\Lambda^{\prime}$, each component of $\Lambda\setminus\{s\}$ is either spherical or affine. Axiomatizing this feature on the diagrams leads to much larger classes of Artin groups which are not necessarily hyperbolic type, but share some common geometric features with the hyperbolic type ones. From this viewpoint, each quasi-Lannér diagram gives rise to one or several families of Artin groups containing examples of arbitrarily high dimension (by contrast, hyperbolic type Artin groups only exist up to a certain dimension), and we will organize the remaining results in this way.
Consideration of cores is part of a strategy to treat all Artin groups; see Section 1.3.
We have treated one case when the core is a cycle in [Hua23]. Theorem 1.3 relies on treating other types of cores, which leads to the following more general result.
@@
(=Proposition 11.1) Given an Artin group $A_{\Lambda}$ with its Dynkin diagram $\Lambda$ containing an induced subdiagram $\Lambda^{\prime}\subset\Lambda$ such that
  1. 1.
    $\Lambda^{\prime}$ is the Dynkin diagram of a 3-dimensional irreducible affine Coxeter group;
  2. 2.
    for any vertex $s\in\Lambda^{\prime}$, each component of $\Lambda\setminus\{s\}$ is either spherical, or has type in $\{\widetilde{A}_{3},\widetilde{B}_{3},\widetilde{C}_{n}\}$.
Then $A_{\Lambda}$ satisfies the $K(\pi,1)$-conjecture.
To demonstrate Theorem 1.4, we give seven families of Artin groups, each family containing examples of arbitrary high dimensions whose $K(\pi,1)$-conjecture was not known before. Some of them can be treated directly by Theorem 1.4, and others require a small variation of Theorem 1.4. We use the following convention for drawing a Dynkin diagram: when an edge is unlabeled, then its label is assumed to be 3.
@@
(=Corollary 11.3) All Artin groups whose Dynkin diagrams belong to Figure 1 satisfy the $K(\pi,1)$-conjecture.
Figure 1: Some variations of quasi-Lannér diagrams
Motivated by a few examples of 5-dimensional quasi-Lannér diagrams (see Figure 2 (I), (II)), we study a family of diagrams, called extended stars, which generalize McCammond’s star diagrams $K_{1,n}$ mentioned before, see Figure 2 (III).
Figure 2: Extended star.

### v3

# Cycles in spherical Deligne complexes and $K(\pi,1)$-conjecture for Artin groups

@@ Abstract
We introduce a method of finding large non-positively curved subcomplexes in certain spherical Deligne complexes, which is effective for studying fillings of certain 6-cycles in spherical Deligne complexes. As applications, we show the $K(\pi,1)$-conjecture holds for all 3-dimensional hyperbolic type Artin groups, except one single example; and the conjecture holds for all quasi-Lannér hyperbolic type Artin groups up to dimension 4. In higher dimension, we show the $K(\pi,1)$-conjecture for Artin groups whose Coxeter diagrams are complete bipartite (edge labels can be arbitrary), answering a question of J. McCammond.
@@ 1 Introduction
1.1 Background and motivation
A Coxeter group $W_{S}$ is a group with generating set $S=\{s_{1},\ldots,s_{n}\}$ and presentation
$\langle s_{1},s_{2},\ldots,s_{n}\mid(s_{i}s_{j})^{m_{ij}}=1\rangle$
with $m_{ii}=2$ for each $i$, $m_{ij}=m_{ji}\geq 2$ is either an integer or $\infty$ for $i\neq j$. Here $m_{ij}=\infty$ means no relation between $s_{i}$ and $s_{j}$. There is an associated Coxeter diagram $\Lambda$ (or Coxeter-Dynkin graph) encoding the non-commutativity of generators of $W_{S}$: the collection of nodes of $\Lambda$ is $S$, and two nodes $s_{i}$ and $s_{j}$ are joined an edge labeled by $m_{ij}$ if $m_{ij}\geq 3$.
For each Coxeter group $W_{S}$, there is an associated Artin group $A_{S}$ (or Artin-Tits group), whose generating set is $S$, and for each pair of generators $s_{i}\neq s_{j}$ with $m_{ij}<\infty$, there is a relator of form $s_{i}s_{j}s_{i}\cdots=s_{j}s_{i}s_{j}\cdots$ with both sides being alternating words of length $m_{ij}$. We will also write $A_{\Lambda}$ and $W_{\Lambda}$ for the Artin group and Coxeter group with Coxeter diagram $\Lambda$.
A reflection of $W_{S}$ is a conjugate of an element in $S$. Let $R$ be the collection of all reflections in $W_{S}$. Recall that $W_{S}$ admits a canonical representation $\rho:W\to GL(n,\mathbb{R})$, such that each element in $R$ acts as a linear reflection on $\mathbb{R}^{n}$, and the action $W_{S}\curvearrowright\mathbb{R}^{n}$ stabilizes an open convex cone $I\subset\mathbb{R}^{n}$, called the Tits cone, where the action of $W_{S}$ is properly discontinuous. For each reflection $r\in W_{S}$, let $H_{r}$ be the set of fixed points of $\rho(r)$ in $I$. The collection of all such $H_{r}$ forms an arrangement of hyperplanes in $I$. Let
@@ A Coxeter group $W_{S}$ is a group with generating set $S=\{s_{1},\ldots,s_{n}\}$ and presentation
The $K(\pi,1)$-conjecture for reflection arrangement complements, due to Arnol’d, Brieskorn, Pham, and Thom, predicts that the space $M(W_{S})$ is aspherical for any Coxeter group $W_{S}$. There is an induced action of $W_{S}$ on $M(W_{S})$, which is free and properly discontinuous. The quotient has its fundamental group isomorphic to the Artin group $A_{S}$.
@@
Historically, the interests in $M(W_{S})$ come from a connection with singularity theory: if $W_{S}$ is of type $A$, $D$, or $E$, then $M(W_{S})/W_{S}$ is the complement of the discriminant of the semi-universal deformation of a simple singularity of the same type [Bri70]. More recently, it is known that many $M(W_{S})$ are closely related to spaces of stability conditions on suitable triangulated categories [Bri09, Ike14, QW18, AW22, Hen22, DHL23, QZ23, HL24], see also the survey article [Hen24]. In a few cases, $M(W_{S})$ coincides with components of certain strata of the moduli space of abelian differential [LM12], and moduli spaces of certain quadratic differential [Qiu24, Appendix C].
The $K(\pi,1)$-conjecture is a major unsettled problem in studying Artin groups. A well-known geometric approach to this conjecture reduces the $K(\pi,1)$-conjecture for any Artin groups to certain properties of a short list of Artin groups, called spherical Artin groups. These are Artin groups associated with finite Coxeter group, and they are completely classified, see the first two columns of Figure 1 for their Coxeter diagrams. The following combines [CD95] and [Bow95].
@@
Suppose for any irreducible spherical Artin groups $A_{S}$, any loop of length $<2\pi$ in the associated spherical Deligne complex $\Delta_{S}$ with Moussong metric has a length non-increasing homotopy to the trivial loop. Then the $K(\pi,1)$-conjecture holds for any Artin group.
Understanding loops of length $<2\pi$ in spherical Deligne complexes also leads to the solution of a highly non-trivial list of other important unsettled questions on more general Artin groups, like word problem, the intersection of parabolic subgroups, computation of symmetrical subgroups as well as normalizers and commensurators of elements, etc [Alt98, Cri00, God07, MW21].
The main difficulty of this approach is to deal with the large space of loops (of length $<2\pi$) in spherical Deligne complexes, especially in high dimensions. To circumvent this difficulty, we proposed another approach in [Hua24a] to $K(\pi,1)$-conjecture, aiming at reducing the $K(\pi,1)$-conjecture for any Artin groups to studying only short loops in the 1-skeleton of the spherical Deligne complexes $\Delta_{S}$. More precisely, we only need to deal with edge loops, which are made of a sequence of edges. The meaning of “short” depends on the type of the associated irreducible spherical Artin groups (see Section 2.6), but they all have length $<2\pi$, and in many cases, we only need to deal with loops with $\leq 6$ edges, which is much smaller compared to the collection of edge loops with length $<2\pi$ in high dimension. This approach in [Hua24a] can be combined with the works mentioned above and [Hae21, Section 8] to answer importance questions on Artin groups other than the $K(\pi,1)$-conjecture, see [Hua24a, Corollary 10.13] for an example.
Despite their important roles, the behavior of these short edge loops is far from being well-understood. In this article, we introduce a new method for studying the geometry of spherical Deligne complexes, which is effective for studying certain short edge loops. We discuss results on short edge loops in Section 1.2, applications to $K(\pi,1)$-conjecture in Section 1.3, and the method for understanding spherical Deligne complexes in Section 1.4.
1.2 Cycles in some spherical Deligne complexes
Each finite Coxeter $W_{S}$ can be obtained by taking an appropriate fundamental domain in the round sphere $\mathbb{S}^{n}$ which is a convex spherical simplex, and considering the group of isometries generated by reflections along co-dimension 1 faces of this simplex. Then $\mathbb{S}^{n}$ is tessellated by copies of this simplex, and such tessellation is preserved by the action of $W_{S}$. The underlying simplicial complex of this tessellation is called the Coxeter complex $\mathfrak{C}_{S}$ of $W_{S}$, which can be described algebraically as follows. Vertices of $\mathfrak{C}_{S}$ are in 1-1 correspondence with left cosets of form $\{gW_{S\setminus\{s\}}\}_{g\in W_{S},s\in S}$, where $W_{S\setminus\{s\}}$ denotes the subgroup of $W$ generated by $S\setminus\{s\}$; a collection of vertices of $\mathfrak{C}_{S}$ span a simplex if the corresponding cosets have non-empty common intersection. Let $A_{S}$ be the associated spherical Artin groups. One can define a similar complex by considering the cosets $\{gA_{S\setminus\{s\}}\}_{g\in A_{S},s\in S}$ instead, which gives the spherical Deligne complex $\Delta_{S}$ associated with the spherical Artin group $A_{S}$. A vertex of $\Delta_{S}$ is of type $\hat{s}$ if it corresponds to $gA_{S\setminus\{s\}}$. The complex $\Delta_{S}$ is a union of infinitely many round spheres, intersecting each other in an intricate pattern.
An $n$-cycle in $\Delta_{S}$ is an edge loop made of $n$-edges. The question we need to understand is that given a “short” $n$-cycle $\omega$ in $\Delta_{S}$, what is a way to fill in $\omega$ in the 2-skeleton of $\Delta_{S}$ by a 2-disk made of the least number of 2-simplices of $\Delta_{S}$. We will assume $\omega$ is embedded and induced, i.e. two non-adjacent vertices in the cycle are also non-adjacent in $\Delta_{S}$, otherwise we are reduced to filling in $m$-cycles with $m<n$. The case $n=3$ follows from [CD95, Lemma 4.3.2], where any such cycle bounds a 2-simplex - this plays a key role in the solution of $K(\pi,1)$-conjecture and word problem for type FC Artin groups. The case $n=4$ follows from [Hua24a, Corollary 8.2], where any such cycle has a center, namely a vertex in $\Delta$ which is adjacent to each vertex of $\omega$, hence $\omega$ is filled by a disk in the 2-skeleton made of four 2-simplices. As most of the cycles we need for $K(\pi,1)$-conjecture satisfy $n\leq 6$ (see Section 2.6), in this article, we focus on the case $n=6$, which is much more subtle compared the cases $n=3$ and $n=4$. It is no longer true that any embedded and induced 6-cycle $\omega$ in $\Delta_{S}$ has a center. Actually, the combinatorics of minimal filling disks of embedded and induced $6$-cycles are highly sensitive to the type of the spherical Artin group $A_{S}$, as well as the types of vertices of $\omega$.
We introduce a geometric method to understand these cycles, see Section 1.4. It is different in nature from the Garside theoretic method [Hua24a] which does not work when $n>4$, and the topological methods in [Cha04] and [Hae21] (based on [BCKM05, Section 2]) which do not work for exceptional Artin groups. This allows us to process cycles which is not accessible by previous methods.
@@ 1.2 Cycles in some spherical Deligne complexes
(=Theorem 5.6+Lemma 12.2) Suppose $A_{S}$ is of type $A_{n},B_{n},H_{3},F_{4}$ with $n\geq 3$. Let $\omega$ be an embedded 6-cycle in $\Delta_{S}$ such that the type of its vertices alternates between $\hat{s}$ and $\hat{t}$, where $\{s,t\}\in S$ does not commute. Then $\omega$ has a center.
The conclusion of this theorem breaks down if $s$ and $t$ commute.
The most interesting cases of Theorem 1.2 are type $A_{n}$ and $H_{3}$. In the exception case $H_{3}$, we prove more. Recall that an induced 6-cycle $\omega$ in $\Delta_{S}$ has quasi-center, if there is a vertex $x\in\Delta_{S}$ such that $x$ is adjacent to three vertices of $\omega$ which are pairwise non-adjacent. Having a quasi-center breaks the 6-cycle $\omega$ into three 4-cycles, hence minimal filling disks of $\omega$ can be understood through minimal filing disks of 4-cycles in [Hua24a].
@@
(=Theorem 7.1) Suppose $A_{S}$ is of type $H_{3}$. Let $\omega$ be an embedded and induced 6-cycle in $\Delta_{S}$. Then $\omega$ has a quasi-center.
Note that the theorem is not true if one replaces “quasi-center” by “center” in the conclusion. It is also interesting to compare to the case when $A_{S}$ is of type $A_{n}$, where generally induced 6-cycles in $\Delta_{S}$ do not have a quasi-center - this can be seen by using Deligne’s description of the spherical Deligne complex in [Del72].
@@
(=Proposition 3.2) Suppose $A_{S}$ is of type $F_{4}$, with $s\in S$ being a terminal node in its Coxeter diagram. Let $\omega$ be an embedded and induced 6-cycle in $\Delta_{S}$ without any vertices of type $\hat{s}$. Then $\omega$ has a quasi-center.
Complementary to the above results, some 6-cycles in type $D_{n}$ case are treated in a companion article [Hua24b], as it uses a different method.
As a comparison to previous work, we record the following intriguing consequence of Theorem 1.2. Recall that for any subset $S^{\prime}\subset S$, the relative Artin complex $\Delta_{S,S^{\prime}}$ is the full subcomplex of $\Delta_{S}$ spanned by vertices of type $\hat{s}$ with $s\in S^{\prime}$.
@@
(=Corollary 5.7) Suppose $A_{S}$ is of type $A_{n}$ with $n\geq 3$. Let $S^{\prime}\subset S$ be a subset made of three consecutive nodes in the Coxeter diagram of $A_{S}$. Then $\Delta_{S,S^{\prime}}$ with its simplices equipped with $A_{3}$-shape is CAT$(1)$.
The case $n=3$ is the main result of [Cha04].
Figure 1: Spherical and affine diagrams: unlabeled edges are assumed to have label 3.
1.3 Application to the $K(\pi,1)$-conjecture
Applications of results in Section 1.2 to $K(\pi,1)$-conjecture divides into two categories, those which are obtained without too much effort, and those that need substantially more work. In this article we discuss the first category. The second category of results will appear elsewhere.
We caution the reader that our results are stated using Coxeter diagrams, which we also called Coxeter-Dynkin diagrams or Dynkin diagrams in [HH23, Hua24a]. There is a different graph used to represent Artin groups in the literature, called the defining graph or presentation graph, which we are not using here.
The $K(\pi,1)$-conjecture for an arbitrary Artin group reduces to the $K(\pi,1)$-conjecture for free-of-infinity Artin groups (i.e. there are no $\infty$-labeled edges in the Coxeter diagram) by work of Charney-Davis, Ellis-Sköldberg and Godelle-Paris [CD95, ES10, GP12]. Among free-of-infinity Artin groups, the $K(\pi,1)$-conjecture is known in the 2-dimension case by Charney-Davis [CD95], and a few families that are not exactly 2-dimensional but have strong 2-dimensional features by work of Juhasz and Goldman [Juh18, Gol24, Juh23].
Known cases for free-of-infinity Artin groups in dimension $\geq 3$ are limited. Charney proved the conjecture for Artin groups with irreducible spherical parabolics being 2-dimensional or $A_{3}$ [Cha04]. The conjecture is known for spherical Artin groups by Deligne [Del72] and affine Artin groups by Paolini-Salvetti [PS21] (based on McCammond-Sulway [MS17]). These two classes correspond to Coxeter diagrams in Figure 1. The conjecture is also proved for Artin groups whose Coxeter diagrams are trees [Hua24a] with some constraints on edge labeling.
As the $K(\pi,1)$ conjecture is settled for Artin groups in Figure 1 whose Coxeter groups acts on $\mathbb{S}^{n}$ and $\mathbb{E}^{n}$ [Del72, PS21], it is natural to ask the same question for hyperbolic type Artin groups, i.e., Artin groups associated with Coxeter groups acting on the $n$-dimensional hyperbolic space $\mathbb{H}^{n}$ with finite covolume. This was completely understood in the case of $\mathbb{H}^{2}$ [CD95], with a recent alternative proof [DPS22]. However, for $\mathbb{H}^{3}$, only a few sporadic examples of Artin groups are known to satisfy the $K(\pi,1)$-conjecture [Cha04, HH23, Hua24a]. As an immediate consequence of Section 1.2, we give an almost complete treatment in the $\mathbb{H}^{3}$ case.